OpenStudy (anonymous):

solve 3x^2-5x-11=0 find solutions and x n y intercepts

OpenStudy (anonymous):

sorry just x intercept

OpenStudy (anonymous):

use quadratic equation to find the solution and, for x intercept, plug in y = 0 for y intercept , plug in x = 0

OpenStudy (anonymous):

can you factor your polynomial?

OpenStudy (anonymous):

as far as i can get with that is (3x-5)(x+2) but i know that is far from right

OpenStudy (anonymous):

what techniques do you know?

OpenStudy (anonymous):

factoring techniques?

OpenStudy (anonymous):

im struggling with it.If it doesnt have an extra variable im ok but when its like this i dont get it

OpenStudy (anonymous):

you can definietly use factoring technique but you will hard time finding its factor

OpenStudy (anonymous):

i hvr a hard time with all period. passed my first math class with no help and this new math class thats all ive done is ask for help.

OpenStudy (anonymous):

yeah I would highly consider taking saruz advice here and computing the quadratic equation here: \[\frac{-b \pm \sqrt{b^{2}-4(a)(c)}}{2a}\]

OpenStudy (anonymous):

3x^2-5x-11=0 where a = 3 b = -5 and c = -11

OpenStudy (anonymous):

thanks for all the helps. as much as i have read and worked problems out I just cant get it. I will probably end up failing agian. Ive got to many problems to work and Its just making it harder on me. thanks though.